OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x^2 + 2x + y^2 + 4y = 20

OpenStudy (anonymous):

Please help :)

OpenStudy (ankit042):

can you convert the equation to (x-h)^2+(y-k)^2= r^2 format?

OpenStudy (ankit042):

also there is a direct formila which you can use but try with the first method first

OpenStudy (anonymous):

No because the 2x is cofusing me

OpenStudy (ankit042):

(a+b)^2 = a^2+b^2+2ab now we know a=x and b is a number. so what should be value of b to make 2ab=2x?

OpenStudy (anonymous):

I have no idea I'm so confused

OpenStudy (ankit042):

lol ok (x+1)^2 can you expand this?

OpenStudy (anonymous):

simplify? expand? I'm so sorry I have no idea what im doing

OpenStudy (ankit042):

expand same way i did for (a+b)^2

OpenStudy (anonymous):

(x+1).(x+1)=?

OpenStudy (anonymous):

x.(x+1)+1(x+1)

OpenStudy (anonymous):

Nevermind thanks anyways I honestly don;t think I can do this

OpenStudy (ankit042):

Nooo don't give up

OpenStudy (ankit042):

(x+1)^2+(y+2)^2 can you expand this?

OpenStudy (anonymous):

I dont understand how to expand it I literally cannot think at all

OpenStudy (ankit042):

hmm you know how to simplify (x+1)^2 ?

OpenStudy (anonymous):

x^2+2?

OpenStudy (anonymous):

What I originally got was x^2+2x)+(y^2+4y)=20...

OpenStudy (ankit042):

it is like muliplying (x+1)(x+1) which is x(x+1)+1(x+1)

OpenStudy (ankit042):

x^2+x+x+1= x^2+2x+1

OpenStudy (ankit042):

so if you remember this formula (a+b)^2 = a^2+b^2+2ab you will be good does any of this make sense?

OpenStudy (anonymous):

Sort of.

OpenStudy (anonymous):

This is my last essay and then I'm done I'm mentally exausted

OpenStudy (ankit042):

haha ok you can try expanding (y+2)^2 lets see else take rest and do it later!

OpenStudy (anonymous):

I can't sadly. I have to turn it in tonight. So oh well

OpenStudy (anonymous):

I did x^2+ax+y^2+by=c then added the variabes as a=2, b=4, and c=20. So x^2+2x+y^2+4y=20. But i didnt know how to finish or if thatwas right so i give up

OpenStudy (ankit042):

hmm it is correct so you can use the formula center = -a/2,-b/2

OpenStudy (anonymous):

so -1, -2?

OpenStudy (ankit042):

and radius is sqrt(c-a^2/4 - b^2/4)

OpenStudy (ankit042):

yeah

OpenStudy (ankit042):

but I will suggest to learn the other way as well, anyway gn :)

OpenStudy (anonymous):

16/-12/4? how do you solve that?

OpenStudy (ankit042):

sqrt(20-1-4)

OpenStudy (anonymous):

2.44?

OpenStudy (anonymous):

oh no 20-12-4

OpenStudy (anonymous):

so 2?

OpenStudy (ankit042):

where is 12 comming from?

OpenStudy (anonymous):

\[x^2 + 2x + y^2 + 4y = 20\]right?

OpenStudy (anonymous):

takes one step what is half of 2 ?